geerky42 (geerky42):

Let f(x) = ax³ - bx + 2 where a and b are constants. If f(4) = -5, then f(-4) = ? A) 9 B) 5 C) 1 D) -1 E) -4

OpenStudy (anonymous):

Do you have an idea about what the answer is?

geerky42 (geerky42):

No idea.

geerky42 (geerky42):

It just seems like there is not enough given information...

OpenStudy (anonymous):

I am guessing that, because if f(4)= -5 that the f(-4)= (-5 x -1) Because 4 x -1 = -4

OpenStudy (anonymous):

if that makes sense

geerky42 (geerky42):

But f(x) is neither even nor odd function...

geerky42 (geerky42):

It doesn't make sense to me.

OpenStudy (anonymous):

I am just ignoring the first part..

OpenStudy (anonymous):

@Mertsj Do you know if I am right?

jimthompson5910 (jim_thompson5910):

plug in x = 4 and y = -5 to get one equation in terms of a and b then plug in x = -4 to get another equation in terms of 'a' and b then add the two equations

jimthompson5910 (jim_thompson5910):

afterwards, solve for y

OpenStudy (anonymous):

Huh? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

y = a*x^3-b*x+2 -5 = a*(4)^3-b*(4)+2 -5 = 64a - 4b + 2 -5 - 2 = 64a - 4b 64a - 4b = -7

jimthompson5910 (jim_thompson5910):

y = a*x^3-b*x+2 y = a*(-4)^3-b*(-4)+2 y = -64a+4b+2 -64a + 4b + 2 = y

jimthompson5910 (jim_thompson5910):

add the two equations 64a - 4b = -7 -64a + 4b + 2 = y ------------------ 0a + 0b + 2 = -7 + y which gives us 2 = -7 + y

geerky42 (geerky42):

Thanks!

OpenStudy (anonymous):

Oh.